Obtain a partial Differential equation by eliminating arbitrary constants from z=(x-α) ^2 +(y-β) ^2?



A) 2z=p^2+q^2



B) 4z=p^2+q^2



C) 4z=p+q



D) z^2=p^2+q^2




1
Expert's answer
2022-02-11T11:12:44-0500

Solution:

"z=(x-α) ^2 +(y-β) ^2\ ...(i)
\\ \Rightarrow \dfrac{\partial z}{\partial x}=2(x-\alpha)
\\ \Rightarrow p=2(x-\alpha)
\\ \Rightarrow (x-\alpha)=\dfrac p2\ ...(ii)"

Next, "\dfrac{\partial z}{\partial y}=2(y-\beta)"

"\Rightarrow q=2(y-\beta)
\\ \Rightarrow (y-\beta)=\dfrac q2 \ ...(iii)"

Put (ii) and (iii) in (i).

"z=(\dfrac p2) ^2 +(\dfrac q2) ^2
\\ \Rightarrow 4z=p^2+q^2"

Option B is correct.